OpenStudy (anonymous):

Describe in paragraph form how to solve this system using elimination. x + 2y = 10 2x +3y = 16

OpenStudy (johnweldon1993):

multiply the top equation by -2 this way the x becomes -2x when you combine the equation you'll see that the -2x and 2x cancel out and then you only have y to solve for

OpenStudy (abb0t):

You could use substitution or addition method. I prefer substitution. Rerrange one of your equations in the form y = ax+n or x = by +n then use that to substitute in for "x" or "y" into your secod equation.

OpenStudy (abb0t):

You should have only one variable in your equation now (only x's or only y's - depending what variable you decided to substitute). Then solve for that variable. and plug in ur value into equation one and solve for the 2nd variable.

OpenStudy (johnweldon1993):

(-2)x + 2y = 10 2x +3y = 16 so it becomes -2x -4y = -20 now you have -2x -4y = -20 and 2x +3y = 16 now add the 2 equations -2x + 2x? = 0 -4y + 3y? -1y or just -y -20 + 16? -4 so you have -y = -4 or y = 4 now subthat 4 in wherever you see y in either equation and solve for x :)

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (johnweldon1993):

anytime :) glad to help :)